# THE RESTRICTED PICARD FUNCTOR

### v2

@@ 4 Descent through purely inseparable extensions
Let $X$ be an affine $K$-scheme, $L/K$ a finite purely inseparable extension field, and $Y:=(X_{L})_{\rm{red}}$. Then ${\rm{DesCan}}_{Y/X}\simeq\operatorname{R}_{L/K}(\operatorname{Pic}^{+}_{Y/L})$.
@@
Because $L/K$ is radicial, $\operatorname{R}_{L/K}$ induces an equivalence of étale sites on $T$ and $T_{L}$ for any $K$-scheme $T$. It follows that $\operatorname{R}_{L/K}$ commutes with étale sheafification, so it suffices to show that one has an isomorphism
${\rm{DesCan}}_{Y/X}^{\rm{naive}}\simeq\operatorname{R}_{L/K}((\operatorname{Pic}^{+}_{Y/L})^{\rm{naive}})$
@@
agree. Because $L/K$ is purely inseparable, the difference between the two projections $L\rightarrow L\otimes_{K}L$ is nilpotent, hence the two projections
${\rm{Spec}}(L\otimes_{K}L)_{\rm{red}}\rightarrow{\rm{Spec}}(L)$
agree. Because $Y$ is a closed subscheme of $X_{L}$, it follows that the same holds for the projections in (4.1). ∎
It remains to verify the local boundedness of the collection of descent data associated to a finite purely inseparable extension $L/K$. The following lemma – akin to an integral domain property for tensor powers of rings – is simple, but lies at the heart of the proof of Theorem 1.3, and in particular, of the boundedness of the collection of descent data for line bundles on a scheme required in Proposition 3.4 in order to descend representability.
@@
Let $B$ be a finite free $A$-algebra, and let $\pi_{12},\pi_{23}\colon B\otimes_{A}B\rightarrow B\otimes_{A}B\otimes_{A}B$ denote the maps $b\otimes b^{\prime}\mapsto b\otimes b^{\prime}\otimes 1$, $1\otimes b\otimes b^{\prime}$, respectively. Let $I,I^{\prime},I^{\prime\prime}$ be ideals of $A$ such that, for $b,b^{\prime}\in B$, if $bb^{\prime}\in I^{\prime\prime}B$, then either $b\in IB$ or $b^{\prime}\in I^{\prime}B$. Then for $\alpha,\beta\in B\otimes_{A}B$, if $\pi_{12}(\alpha)\pi_{23}(\beta)\in I^{\prime\prime}(B\otimes_{A}B\otimes_{A}B)$, then either $\alpha\in I(B\otimes_{A}B)$ or $\beta\in I^{\prime}(B\otimes_{A}B)$.
@@
We shall also require the following lemma about flatness of reductions of base changes to larger fields.
Lemma 4.3.
[…]
Proof.
We may assume that $X={\rm{Spec}}(R)$ with $R$ a normal domain. Let $F:={\rm{Frac}}(R)$, and view $F$ and $L$ as lying in a fixed algebraic closure of $F$. We will prove:
  • (1)
    If $L\subset F$, then the composition $R\rightarrow R_{L}\rightarrow(R_{L})_{\rm{red}}$ is an isomorphism.
  • (2)
    If $L\cap F=K$, then $R_{L}$ is reduced.
These two assertions suffice to establish the lemma, as - because $[L:K]=p$ we must be in either situation (1) or (2).
First let us prove (1). The map $R\rightarrow(R_{L})_{\rm{red}}$ is injective because $R$ injects into $R_{L}$. The key point is therefore surjectivity. We first note that surjectivity holds in the case $R=L$. Indeed, $(L\otimes_{K}L)_{\rm{red}}$ is a finite, reduced, radicial extension of $L$, hence a finite extension field $E/L$. But the map $L\rightarrow L\otimes_{K}L$ admits a section, namely, the multiplication map, hence $L\rightarrow E$ also admits a section, so $E=L$. For general $R$, because $R$ is a normal $K$-algebra and $L\subset F$ is algebraic over $K$, one has $L\subset R$. To prove surjectivity of $R\rightarrow(R\otimes_{K}L)_{\rm{red}}$, we only need show that anything of the form $1\otimes\ell$ with $\ell\in L$ lies in the image, and in fact any such element lies in the image of $L\subset R$ because of the already-treated case $R=L$. This proves (1).
[…]
The following somewhat technical lemma will be used in conjunction with Proposition 4.2 to prove the boundedness of the descent data arising from descent through a finite purely inseparable extension.
@@
We may now verify the local boundedness of the descent data for a purely inseparable extension of degree $p$. For the notion of local boundedness of a family of isomorphisms, see Definition 3.1.
@@
Let $L/K$ be a purely inseparable, degree $p$ extension of fields, $X$ a normal separated $K$-scheme of finite type, and $Y:=(X_{L})_{\rm{red}}$. Let $T$ be a smooth $K$-scheme, and $\mathscr{L}$ a line bundle on $Y_{T}$ such that $\pi_{1}^{*}(\mathscr{L})\simeq\pi_{2}^{*}(\mathscr{L})\in\operatorname{Pic}((Y\times_{X}Y)_{T})$. Then the collection of all descent data for $\mathscr{L}_{S}$ with respect to the fppf cover $Y_{S}\rightarrow X_{S}$ is locally bounded as $S$ varies over all $T$-schemes smooth over $K$.
@@
We may extend scalars and assume that $K$ is separably closed. Further, we may work locally on $T$ and thereby assume that $T$ is quasi-compact, and in this case we must prove boundedness of the collection of descent data. Because the $K$-points of $S$ are Zariski dense in $S$ for any $K$-smooth $S$, we may assume that $S={\rm{Spec}}(K)$. To prove boundedness of the descent data, recall that a descent datum is an isomorphism between the pullbacks of $\mathscr{L}$ along the two projections $Y\times_{X}Y\rightrightarrows Y$ which additionally satisfies the cocycle condition. Fix an isomorphism $\phi\colon\pi_{1}^{*}(\mathscr{L})\xrightarrow{\sim}\pi_{2}^{*}(\mathscr{L})$. Let $s\colon{\rm{Spec}}(K)\rightarrow T$ be the $T$-scheme structure on $S={\rm{Spec}}(K)$. Let $\pi_{ij}\colon Y\times_{X}Y\times_{X}Y\rightarrow Y\times_{X}Y$ denote the various projections. Then the cocycle condition is equivalent to the equation
@@
where $C:=s^{*}(C^{\prime})$ with $C^{\prime}:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi)^{-1}\pi_{12}^{*}(\phi)^{-1}\in\Gamma((Y\times_{X}Y\times_{X}Y)\times_{K}T,{\mathbf{G}}_{\rm{m}})$. Because $T$ is quasi-compact, there is a finite-dimensional $W\subset\Gamma((Y^{\times_{X}3}),\mathcal{O})$ such that $C^{\prime},(C^{\prime})^{-1}\in W_{T}$. Now we forget the some of the particulars of our underlying situation, and we simply prove the following: For any finite-dimensional $W\subset\Gamma(Y^{\times_{X}3},\mathcal{O})$, there is a finite-dimensional $V\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ such that any $u\in\Gamma(Y\times_{X}Y,\mathcal{O})$ satisfying
@@
lies in $V$. Then we see that all of the $(t,u)$ defining a cocycle satisfy $u,u^{-1}\in V$.
To prove the claim, we first choose a normal compactification $\overline{X}$ of $X$, and let $\overline{Y}:=(\overline{X}_{L})_{\rm{red}}$, which is a compactification of $Y$ living in a commutative diagram
@@
with $\overline{f}$ radicial and flat by Lemma 4.3. Let $D_{1},\dots,D_{n}$ be the boundary divisors of $\overline{X}$ (that is, those divisors not contained in $X$), and let $D:=\sum_{i=1}^{n}D_{i}$. For each $N>0$, consider the coherent sheaf $\mathcal{O}_{\overline{X}}(ND)$. Let $g\colon\overline{Y}\times_{\overline{X}}\overline{Y}\rightarrow\overline{X}$ be the natural map, and consider the open embedding $j\colon Y\times_{X}Y\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$. Because the support of $g^{-1}(ND)$ is disjoint from $Y\times_{X}Y$, one has a canonical isomorphism $j^{*}g^{*}(\mathcal{O}_{\overline{X}}(ND))\xrightarrow{\sim}\mathcal{O}_{Y\times_{X}Y}$ which is adjoint to a map $g^{*}(\mathcal{O}_{\overline{X}}(ND))\rightarrow j_{*}(\mathcal{O}_{Y\times_{X}Y})$. Let $\mathscr{G}_{N}$ denote the image of this map, a coherent sheaf on $\overline{Y}\times_{\overline{X}}\overline{Y}$. We will show that, for some $N>0$ depending on $W$, every $u$ satisfying (4.2) lies in ${\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},\mathscr{G}_{N})\subset{\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},j_{*}\mathcal{O}_{Y\times_{X}Y})={\rm{H}}^{0}(Y\times_{X}Y,\mathcal{O}_{Y\times_{X}Y})$. Because $\overline{Y}\times_{\overline{X}}\overline{Y}$ is proper over $K$ and $\mathscr{G}_{N}$ is coherent, this will show that the set of such $u$ lies in a finite-dimensional $K$-vector space.
To begin with, we claim that being in $\mathscr{G}_{N}$ is determined by codimension-one points. More precisely, let $Z=\coprod{\rm{Spec}}(\mathcal{O}_{s})$ be the disjoint union of the (spectra of the) local rings of all codimension-one points of $\overline{Y}\times_{\overline{X}}\overline{Y}$, and let $h\colon Z\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$ be the natural map. Then we claim that a section of ($j_{*}$ of) the structure sheaf of $Y\times_{X}Y$ lies in $\mathscr{G}_{N}$ if and only if the same holds upon pulling back to $Z$. This will follow if we can show that the adjunction map $g^{*}(\mathcal{O}(ND))\rightarrow h_{*}h^{*}(g^{*}(\mathcal{O}(ND)))$ is an isomorphism, and similarly with $g^{*}(\mathcal{O}(ND))$ replaced by $\mathcal{O}_{\overline{Y}\times_{\overline{X}}\overline{Y}}$. To check this, note that the analogous statement holds for $\mathcal{O}_{\overline{X}}(ND)$ and the structure sheaf on $\overline{X}$ because $\overline{X}$ is normal: If we define $R$ to be the disjoint union of the spectra of the local rings of $\overline{X}$ at codimension-one points, and $r\colon R\hookrightarrow\overline{X}$ the inclusion, then the corresponding adjunction map is an isomorphism for $\mathcal{O}_{\overline{X}}(ND)$ and for $\mathcal{O}_{\overline{X}}$. Now we claim that if a quasi-coherent sheaf $\mathscr{F}$ on $\overline{X}$ is such that $\mathscr{F}\rightarrow r_{*}r^{*}(\mathscr{F})$ is an isomorphism, then the same holds for the map $g^{*}(\mathscr{F})\rightarrow h_{*}h^{*}(g^{*}(\mathscr{F}))$. To see this, consider the following commutative diagram, which is even Cartesian because $g$ is a homeomorphism:
@@
If the map $\mathscr{F}\rightarrow r_{*}r^{*}(\mathscr{F})$ is an isomorphism, then so is $g^{*}(\mathscr{F})\rightarrow g^{*}r_{*}r^{*}(\mathscr{F})$. Because $\overline{f}$ is flat, so is $g$, hence pullback along $g$ commutes with pushforward of quasi-coherent sheaves. Thus $g^{*}r_{*}r^{*}(\mathscr{F})=h_{*}t^{*}r^{*}(\mathscr{F})=h_{*}h^{*}(g^{*}(\mathscr{F}))$. So the map $g^{*}(\mathscr{F})\rightarrow h_{*}h^{*}(g^{*}(\mathscr{F}))$ is also an isomorphism. This proves the claim that whether a section of $j_{*}(\mathcal{O}_{Y\times_{X}Y})$ lies in $\mathscr{G}_{N}$ may be checked at codimension-one points.
@@
Note that for any $u\in{\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},j_{*}\mathcal{O}_{Y\times_{X}Y})$, $u|Y\times_{X}Y$ lives in $g^{*}(\mathcal{O}(ND))|Y\times_{X}Y=\mathcal{O}$. Thus the only issue is whether $u$ lies in the image of $g^{*}(\mathcal{O}(ND))$ at the codimension-one points of the boundary of $\overline{Y}\times_{\overline{X}}\overline{Y}$. Because $\overline{Y}\times_{\overline{X}}\overline{Y}$ is a closed subscheme of $\overline{X}\otimes_{K}(L\otimes_{K}L)$, sections of the structure sheaf of $\overline{Y}\times_{\overline{X}}\overline{Y}$ all lie in the $(L\otimes_{K}L)$-span of sections of $\mathcal{O}_{\overline{X}}$, and for any section s of $\mathcal{O}_{\overline{X}}$, one has that $s\in\mathcal{O}_{\overline{X}}(nD)$ for some $n>0$ depending on $s$. It follows that any section of $\mathcal{O}_{\overline{Y}\times_{\overline{X}}\overline{Y}}$ lies in $\mathscr{G}_{n}$ for some $n>0$. We must show that there is a uniform bound on the $n$ required as $u$ varies over sections satisfying (4.2). Because $W$ is finite-dimensional, one has $W\subset\mathscr{G}_{m}$ for some $m>0$.
@@
Let $z$ be one the finitely many boundary points of $\overline{Y}\times_{\overline{X}}\overline{Y}$. We will show that there is $N>0$ depending on $z$ such that every $u$ satisfying (4.2) lies in $\mathscr{G}_{N}$ near $z$. Since there are only finitely many $z$, this will suffice. Let $x:=g(z)$ be the image of $z$ in $\overline{X}$, and let $y\in\overline{Y}$ be the (unique) boundary point lying above $x$. Also let $B:=\mathcal{O}_{\overline{Y},y}$, $A:=\mathcal{O}_{\overline{X},x}$. Then $A$ is a Japanese – excellent even – DVR, and $B$ is finite flat over $A$. Furthermore, because $Y=(X_{L})_{\rm{red}}$, $B$ is spanned over $A$ by $L$. Thus $B^{p}\subset A$. We are therefore in the situation of Lemma 4.4. We claim that we may take $N=m+r-1$, where $r$ is as in that lemma.

### v4

@@ 4 Descent through radicial covers
Let $X$ be an irreducible affine $K$-scheme, and let $f\colon Y\rightarrow X$ be an affine radicial map. Then ${\rm{DesCan}}_{Y/X}\simeq\operatorname{Pic}^{+}_{Y/K}$.
@@
It suffices to show that one has an isomorphism
${\rm{DesCan}}_{Y/X}^{\rm{naive}}\simeq(\operatorname{Pic}^{+}_{Y/K})^{\rm{naive}}$
@@
agree. Call these two maps $g_{1},g_{2}$. Let $j\colon\eta\rightarrow((Y\times_{K}Y)_{T})_{\rm{red}}$ be the inclusion of one of the generic points of $((Y\times_{K}Y)_{T})_{\rm{red}}$. It suffices to show that the maps $g_{i}\circ j$ agree. But the $g_{i}\circ j$ agree after postcomposing with $f_{T}\colon Y_{T}\rightarrow X_{T}$. Since $f_{T}$ is radicial, it follows that the $g_{i}\circ j$ agree. ∎
Next we prove the local boundedness of the collection of descent data associated to a radicial extension. The following lemma – akin to an integral domain property for tensor powers of rings – is simple, but lies at the heart of the proof of Theorem 1.3, and in particular, of the boundedness of the collection of descent data for line bundles on a scheme required in Proposition 3.4 in order to descend representability.
@@
Let $A$ be a ring, let $B$ be a finite free $A$-algebra, and let $\pi_{12},\pi_{23}\colon B\otimes_{A}B\rightarrow B\otimes_{A}B\otimes_{A}B$ denote the maps $b\otimes b^{\prime}\mapsto b\otimes b^{\prime}\otimes 1$, $1\otimes b\otimes b^{\prime}$, respectively. Let $I,I^{\prime},I^{\prime\prime}$ be ideals of $A$ such that, for $b,b^{\prime}\in B$, if $bb^{\prime}\in I^{\prime\prime}B$, then either $b\in IB$ or $b^{\prime}\in I^{\prime}B$. Then for $\alpha,\beta\in B\otimes_{A}B$, if $\pi_{12}(\alpha)\pi_{23}(\beta)\in I^{\prime\prime}(B\otimes_{A}B\otimes_{A}B)$, then either $\alpha\in I(B\otimes_{A}B)$ or $\beta\in I^{\prime}(B\otimes_{A}B)$.
@@
The following somewhat technical lemma will be used in conjunction with Proposition 4.2 to prove the boundedness of the descent data arising from descent through a radicial cover.
@@
We may now verify the local boundedness of the descent data associated to a radicial cover. For the notion of local boundedness of a family of isomorphisms, see Definition 3.1.
@@
Let $f\colon Y\rightarrow X$ be a finite flat radicial morphism between connected normal $K$-schemes of finite type. Let $T$ be a smooth $K$-scheme, and $\mathscr{L}$ a line bundle on $Y_{T}$ such that $\pi_{1}^{*}(\mathscr{L})\simeq\pi_{2}^{*}(\mathscr{L})\in\operatorname{Pic}((Y\times_{X}Y)_{T})$. Then the collection of all descent data for $\mathscr{L}_{S}$ with respect to the fppf cover $Y_{S}\rightarrow X_{S}$ is locally bounded as $S$ varies over all $T$-schemes smooth over $K$.
@@
We may assume that $Y\neq\emptyset$. The function field extension $K(Y)/K(X)$ is purely inseparable, so when ${\rm{char}}(K)=0$ it is an isomorphism. Since $f$ is finite and $X$ and $Y$ are normal, it follows that $f$ is an isomorphism, so the proposition is immediate in this case. So from now on we assume that ${\rm{char}}(K)=p>0$. We may extend scalars and assume that $K$ is separably closed. Further, we may work locally on $T$ and thereby assume that $T$ is quasi-compact, and in this case we must prove boundedness of the collection of descent data. Because the $K$-points of $S$ are Zariski dense in $S$ for any $K$-smooth $S$, we may assume that $S={\rm{Spec}}(K)$. To prove boundedness of the descent data, recall that a descent datum is an isomorphism between the pullbacks of $\mathscr{L}$ along the two projections $Y\times_{X}Y\rightrightarrows Y$ which additionally satisfies the cocycle condition. Fix an isomorphism $\phi\colon\pi_{1}^{*}(\mathscr{L})\xrightarrow{\sim}\pi_{2}^{*}(\mathscr{L})$. Let $s\colon{\rm{Spec}}(K)\rightarrow T$ be the $T$-scheme structure on $S={\rm{Spec}}(K)$. Let $\pi_{ij}\colon Y\times_{X}Y\times_{X}Y\rightarrow Y\times_{X}Y$ denote the various projections. Then the cocycle condition on $u\cdot s^{*}(\phi)$ with $u\in\Gamma((Y\times_{X}Y)_{T},{\mathbf{G}}_{\rm{m}})$ is equivalent to the equation
@@
where $C:=s^{*}(C^{\prime})$ with $C^{\prime}:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi)^{-1}\pi_{12}^{*}(\phi)^{-1}\in\Gamma((Y\times_{X}Y\times_{X}Y)\times_{K}T,{\mathbf{G}}_{\rm{m}})$. Because $T$ is quasi-compact, there is a finite-dimensional $W\subset\Gamma(Y\times_{X}Y\times_{X}Y,\mathcal{O})$ such that $C^{\prime},(C^{\prime})^{-1}\in W_{T}$. Now we forget some of the particulars of our underlying situation, and we simply prove the following: For any finite-dimensional $W\subset\Gamma(Y\times_{X}Y\times_{X}Y,\mathcal{O})$, there is a finite-dimensional $V\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ such that any $u\in\Gamma(Y\times_{X}Y,\mathcal{O})$ satisfying
@@
lies in $V$. Then we see that all of the pairs $(s,u)$ as above defining a cocycle satisfy $u,u^{-1}\in V$.
To prove the claim, we first choose a normal compactification $\overline{X}$ of $X$, and let $\overline{Y}$ be the normal closure of $\overline{X}$ in the function field of $Y$. Then $\overline{Y}$ is a compactification of $Y$ living in a commutative diagram
@@
We claim that $\overline{f}$ is radicial, and that the open locus $U\subset\overline{X}$ of points above which $\overline{f}$ is flat is a dense open with ${\rm{codim}}_{\overline{X}}(\overline{X}-U)\geq 2$. First, to see that $\overline{f}$ is radicial, let $V:={\rm{Spec}}(R)\subset\overline{X}$ be affine, and let ${\rm{Spec}}(S)=\overline{f}^{-1}(V)$. Let $t>0$ such that $K(Y)^{p^{t}}\subset K(X)$. Then for any $s\in S$, $s^{p^{t}}\in K(X)={\rm{Frac}}(R)$ is integral over $R$. Because $\overline{X}$ is normal, therefore, $s^{p^{t}}\in R$. Thus,
$\overline{f}_{*}\mathcal{O}_{\overline{Y}}^{p^{t}}\subset\mathcal{O}_{\overline{X}}.$ (4.3)
It follows that ${\rm{Spec}}(S)\rightarrow{\rm{Spec}}(R)$ is radicial. To prove the flatness claim, it suffices to prove that $\overline{f}$ is flat above each codimension $1$ point of $x\in\overline{X}$. Since $\overline{X}$ is normal, the local ring of $x$ is a DVR. Thus we seek to prove that the local ring of $\overline{Y}$ at any point above $x$ is $\mathcal{O}_{\overline{X},x}$-torsion free, and this follows from the fact that $\overline{f}$ is dominant and $\overline{Y}$ is integral.
Let $D_{1},\dots,D_{n}$ be the boundary divisors of $\overline{X}$ (that is, those irreducible divisors not intersecting $X$), and let $D:=\sum_{i=1}^{n}D_{i}$. For each $N>0$, consider the coherent sheaf $\mathcal{O}_{\overline{X}}(ND)$. Let $g\colon\overline{Y}\times_{\overline{X}}\overline{Y}\rightarrow\overline{X}$ be the natural map, and let $V:=g^{-1}(U)\subset\overline{Y}\times_{\overline{X}}\overline{Y}$. Note that $Y\times_{X}Y\subset V$ because $Y$ is flat over $X$. Consider the open embeddings $j\colon Y\times_{X}Y\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$ and $j_{V}\colon V\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$. Because the restriction of $g$ to $V\rightarrow U$ is flat by our choice of $U$, we have
${\rm{Ass}}_{V}(j_{V}^{*}g^{*}\mathcal{O}_{\overline{X}}(ND))\subset\overline{g}^{-1}({\rm{Ass}}_{U}(\mathcal{O}_{\overline{X}}(ND)))=\eta,$
where $\eta$ is the unique point above the generic point of $U$. By [EGA, IV2, Cor. 5.11.4], therefore, the sheaf
$\mathscr{H}_{N}:=(j_{V})_{*}j_{V}^{*}g^{*}\mathcal{O}_{\overline{X}}(ND)$
@@
on $\overline{Y}\times_{\overline{X}}\overline{Y}$ is coherent. Because $Y\times_{X}Y\subset V$, we have $\mathscr{H}_{N}|_{Y\times_{X}Y}\simeq g^{*}\mathcal{O}(ND)|_{Y\times_{X}Y}\simeq\mathcal{O}_{Y\times_{X}Y}$, the last isomorphism because $Y$ lies over $X$, and $\mathcal{O}(ND)|_{X}\simeq\mathcal{O}_{X}$. We thus have an isomorphism $j^{*}\mathscr{H}_{N}\xrightarrow{\sim}\mathcal{O}_{Y\times_{X}Y}$ which is adjoint to a map $\mathscr{H}_{N}\rightarrow j_{*}\mathcal{O}_{Y\times_{X}Y}$. Let $\mathscr{G}_{N}$ be the image of this map. It is a coherent sheaf, hence ${\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},\mathscr{G}_{N})$ is a finite-dimensional $K$-vector space. We will show that, for a suitable $N$, every $u$ satisfying (4.2) lies in ${\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},\mathscr{G}_{N})\subset{\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},j_{*}\mathcal{O})={\rm{H}}^{0}(Y\times_{X}Y,\mathcal{O})$.
Now we claim that membership in $\mathscr{G}_{N}$ is determined by codimension-one points on the boundary of $\overline{Y}\times_{\overline{X}}\overline{Y}$. More precisely, we have the following lemma.
Lemma 4.5.
Let $Z$ be the disjoint union of the spectra of the local rings of $\overline{Y}\times_{\overline{X}}\overline{Y}$ at the generic points of $g^{-1}(D_{1}),\dots,g^{-1}(D_{n})$, and let $h\colon Z\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$ be the natural map. Let $\mathscr{C}$ be the cokernel of the map $\mathscr{G}_{N}\rightarrow j_{*}\mathcal{O}_{Y\times_{X}Y}$. Then the natural map $\mathscr{C}\rightarrow h_{*}h^{*}\mathscr{C}$ is an inclusion.
Proof.
Consider the following diagram, in which the horizontal arrows are the inclusions and the vertical arrows are induced by $g$:
@@
The isomorphism $j_{XU}^{*}\mathcal{O}_{U}(ND)\xrightarrow{\sim}\mathcal{O}_{X}$ is adjoint to an injective map $\mathcal{O}_{U}(ND)\rightarrow(j_{XU})_{*}\mathcal{O}_{X}$. Let $\mathscr{A}$ be the cokernel of this map, so we have an exact sequence of quasi-coherent sheaves
$0\longrightarrow\mathcal{O}_{U}(ND)\longrightarrow(j_{XU})_{*}\mathcal{O}_{X}\longrightarrow\mathscr{A}\longrightarrow 0$ (4.4)
Let $R$ be the disjoint union of the spectra of the local rings of $\overline{X}$ at the generic points of $D_{1},\dots,D_{n}$, and let $r\colon R\hookrightarrow U$ be the inclusion. (Note that $R$ does indeed map into $U$ because ${\rm{codim}}_{\overline{X}}(\overline{X}-U)\geq 2$.) Because $U\subset\overline{X}$ is normal, the canonical map
$\mathscr{A}\rightarrow r_{*}r^{*}\mathscr{A}$ (4.5)
is an inclusion. Let $h^{\prime}\colon Z\rightarrow V$ be the inclusion (so $h=j_{V}\circ h^{\prime}$). Consider the following commutative diagram, which is Cartesian because $g$ is radicial:
[…]
Because $g_{U}$ is flat and $r$ is qcqs, the base change map $g_{U}^{*}r_{*}r^{*}\mathscr{A}\rightarrow h^{\prime}_{*}t^{*}r^{*}\mathscr{A}=h^{\prime}_{*}(h^{\prime})^{*}g_{U}^{*}\mathscr{A}$ is an isomorphism. Thus, applying $g_{U}^{*}$ to the inclusion (4.5) yields that the canonical adjunction map
$g_{U}^{*}\mathscr{A}\hookrightarrow h^{\prime}_{*}(h^{\prime})^{*}g_{U}^{*}\mathscr{A}$
is an inclusion. Then applying $(j_{V})_{*}$ and using the equality $(j_{V})_{*}(h^{\prime})_{*}(h^{\prime})^{*}=h_{*}h^{*}(j_{V})_{*}$, we obtain that the adjunction
$(j_{V})_{*}g_{U}^{*}\mathscr{A}\hookrightarrow h_{*}h^{*}(j_{V})_{*}g_{U}^{*}\mathscr{A}$ (4.6)
is an inclusion.
Since $g_{U}$ is flat, applying $g_{U}^{*}$ to (4.4) yields an exact sequence
$0\longrightarrow g_{U}^{*}\mathcal{O}_{U}(ND)\longrightarrow g_{U}^{*}(j_{XU})_{*}\mathcal{O}_{X}\longrightarrow g_{U}^{*}\mathscr{A}\longrightarrow 0$ (4.7)
Because $g_{U}$ is flat and $j_{XU}$ is qcqs, we have an equality $g_{U}^{*}(j_{XU})_{*}\mathcal{O}_{X}\simeq(j_{Y\times_{X}Y,V})_{*}g_{X}^{*}\mathcal{O}_{X}\simeq(j_{Y\times_{X}Y,V})_{*}\mathcal{O}_{Y\times_{X}Y}$. Substituting this into (4.7) and applying $(j_{V})_{*}$ then yields the exact sequence
$\mathscr{G}_{N}\longrightarrow j_{*}\mathcal{O}_{Y\times_{X}Y}\longrightarrow(j_{V})_{*}g_{U}^{*}\mathscr{A}.$
We therefore see that $\mathscr{C}$ is a subsheaf of $(j_{V})_{*}g_{U}^{*}\mathscr{A}$, so the lemma follows from the inclusion (4.6). ∎
@@
For any section s of $\mathcal{O}_{\overline{X}}$, one has that $s\in\mathcal{O}_{\overline{X}}(nD)$ for some $n>0$ depending on $s$. It follows that any section $u$ of $j_{*}\mathcal{O}_{Y\times_{X}Y}$ lies in $\mathscr{H}_{n}$ for some $n>0$. We must show that there is a uniform bound on the $n$ required as $u$ varies over global sections satisfying (4.2). Because $W$ is finite-dimensional, one similarly concludes that $W|_{\epsilon^{-1}(U)}\subset\epsilon^{*}(\mathcal{O}_{\overline{X}}(mD))$ for some $m>0$, where $\epsilon\colon\overline{Y}\times_{\overline{X}}\overline{Y}\times_{\overline{X}}\overline{Y}\rightarrow\overline{X}$ is the natural map.
@@
Let $z$ be one the finitely many codimension-$1$ boundary points of $\overline{Y}\times_{\overline{X}}\overline{Y}$. We will show that there is $N>0$ depending on $z$ such that every $u$ satisfying (4.2) lies in $\mathscr{G}_{N}$ near $z$. Since there are only finitely many $z$, this will suffice. Let $x:=g(z)$ be the image of $z$ in $\overline{X}$, and let $y\in\overline{Y}$ be the (unique) boundary point lying above $x$. Also let $B:=\mathcal{O}_{\overline{Y},y}$, $A:=\mathcal{O}_{\overline{X},x}$. Then $A$ is a Japanese – excellent even – DVR, and $B$ is finite flat over $A$. Furthermore, by 4.3, $B^{p^{t}}\subset A$. We are therefore in the situation of Lemma 4.3. We claim that we may take $N=m+r-1$, where $r$ is as in that lemma.